OpenStudy (anonymous):

Find the next three terms of the sequence. Then write a rule for the sequence. 648, 216, 72, 24

OpenStudy (anonymous):

8, 8/3, 8/9 It just divides by three every time

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

648/3=216 216/3=72 72/3=24 24/3=8 8/3= 2.6 2.6/3=.888 in terms of decimal instead of fractions its 8, 2.6, 0.88

OpenStudy (anonymous):

so i just write 8, 2.6,0.88?

OpenStudy (amistre64):

the sequence seems to be a geometric progression ....

OpenStudy (amistre64):

spose you divide all the numbers there by say 24; what numbers would we get as a result?

OpenStudy (anonymous):

648/ 24?

OpenStudy (amistre64):

yeah, that should at least get the sequence of numbers to something more readily seen

OpenStudy (anonymous):

what should i write down .. just the number or how he should 648/3....

OpenStudy (anonymous):

showed*

OpenStudy (amistre64):

i get a simpler sequence as: (648, 216, 72, 24,...)/24 = (27,9,3,1, ...) 27,9,3,1,.... : these numbers follow the same pattern displayed but are easier to work with and notice a working pattern

OpenStudy (amistre64):

each new number is 1/3 the previous number; which is the rule for generating the sequence

OpenStudy (anonymous):

thats the question tho... =(

OpenStudy (amistre64):

working with the simpler set of numbers; do you see a pattern developing? doesnt the setup resemble the 3s times table in reverse?

OpenStudy (amistre64):

hmm, prolly not the times table ... but like 3^x seems more suited to it

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

since 3^3 = 27 3^2 = 9 3^1 = 3 3^0 = 1 lets write this as \[3^3,3^2,3^1,3^0,...\] would you agree that a pattern is forming? say the next term is \(3^{-1}\) followed by say \(3^{-2}\) .....

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

given a previous term \(a_{n-1}\) ;each new term \(a_n\) is formed by dividing the previous term by 3: \[a_n=\frac13a_{n-1}\]

OpenStudy (amistre64):

this is the "rule" that defines the sequence